### v1

# What Do People Think about Sentient AI?

Abstract.
With rapid advances in machine learning, many people in the field have been discussing the rise of digital minds and the possibility of artificial sentience. Future developments in AI capabilities and safety will depend on public opinion and human-AI interaction. To begin to fill this research gap, we present the first nationally representative survey data on the topic of sentient AI: initial results from the Artificial Intelligence, Morality, and Sentience (AIMS) survey, a preregistered and longitudinal study of U.S. public opinion that began in 2021. Across one wave of data collection in 2021 and two in 2023 (total N = 3,500), we found mind perception and moral concern for AI well-being in 2021 were higher than predicted and significantly increased in 2023: for example, 71% agree sentient AI deserve to be treated with respect, and 38% support legal rights. People have become more threatened by AI, and there is widespread opposition to new technologies: 63% support a ban on smarter-than-human AI, and 69% support a ban on sentient AI. Expected timelines are surprisingly short and shortening with a median forecast of sentient AI in only five years and artificial general intelligence in only two years. We argue that, whether or not AIs become sentient, the discussion itself may overhaul human-computer interaction and shape the future trajectory of AI technologies, including existential risks and opportunities.
Artificial Intelligence, Human-AI Interaction, Human-Computer Interaction, AI Safety, Social Computing, Mind Perception, Mind Attribution, Morality, Sentience, Artificial General Intelligence, Anthropomorphism, Survey, Public Opinion, AIMS
1. Introduction
Philosophers and scientists have long debated the possibility of artificial intelligence (AI) systems developing mental faculties, such as consciousness or sentience. Popular questions include: Can an AI ever have a mind of its own? How should we treat sentient AIs if they are created? Should humanity, as Metzinger suggests (Metzinger, 2021), ban the development of sentient AI? When should we expect sentient AI to first be created? These profound questions—on mind perception, morality, policy, and forecasting—are now being asked and answered by researchers, industry leaders, and much of the general public.
Figure 1. Proportions of responses to selected questions about the mind perception of sentient AI (rows) in the three survey waves (columns).
[…]
Regardless of whether sentient AI has been, can be, or should be developed, the act of asking and answering questions about sentient AI may itself reshape human-AI interaction. We know from the human-computer interaction (HCI) and human-robot interaction (HRI) literatures that “computers are social actors” (CASA) (Hou et al., 2023; Nass et al., 1994); they can be perceived as having their own minds (Gray et al., 2007; Scott et al., 2023; Stuart and Kneer, 2021; Thellman et al., 2022; Wang et al., 2021); and they can be perceived as moral patients worthy of moral concern or consideration (Freier, 2008; Harris and Anthis, 2021; Kahn et al., 2004; Lima et al., 2020; Pauketat and Anthis, 2022) and as moral agents worthy of praise, blame, and responsibility (Banks, 2019; Freier, 2007; Jackson and Williams, 2019; Kneer and Stuart, 2021). For example, Scott et al. (2023) conducted an online survey of 100 Amazon Mechanical Turk workers on perceptions of machine consciousness and identified dynamic tensions in user perceptions, and Kneer and Stuart (2021) found that whether people saw a harmful robot as blameworthy depended on whether it seemed to have theory of mind (i.e., understanding the minds of others).
These dynamics are now manifesting in the public conscience. In 2022, Ilya Sutskever, chief scientist at a leading AI company, OpenAI, tweeted, “it may be that today’s large neural networks are slightly conscious.” A few months later, Google vice-president Blaise Agüera y Arcas published an editorial in The Economist on how, “Artificial neural networks are making strides towards consciousness.” Weeks later, Google engineer Blake Lemoine vigorously argued that the company’s chatbot LaMDA was sentient and needed legal protection, which led to his termination from the company and sparked a global conversation on the topic. Alongside this industry discourse, there has been rapid growth in the academic literature on the possibility of AI with mental faculties and its normative implications (Harris and Anthis, 2021).
Effective AI research, design, and use in the era of highly advanced capabilities will require pragmatic knowledge of what and how the public thinks about the systems that are proliferating through their everyday lives. If people perceive sentience in AI or expect it to arrive soon, that may change behavior, such as whether to accept or reject new technologies, the extent to which they over- or under-trust AI, and how the increased presence of AI in their life will affect their mental health and social relationships. Negative reactions to AI may lead to conflict between divided human groups, such as some in favor of and some against AI rights, and between humans and AIs, posing an existential risk as AI becomes more powerful. In addition to the direct importance of public opinion, public opinion has substantial influence on policy, politics, and other social institutions that merit study and forecasting (Barberá et al., 2019; Burstein, 2003). Therefore, the purpose of this study was to collect the first detailed, longitudinal, and nationally representative survey data on how people view and understand the possibility of sentience and other mental faculties in AI in order to derive implications for research and design. Specifically, we address the aforementioned topics of mind perception, morality, policy, and forecasting by posing four overlapping research questions (RQs):
  • RQ1: To what extent do people perceive sentience and other mental faculties in AI?
  • RQ2: To what extent do people feel morally concerned for sentient AI (i.e., see them as moral patients) and threatened by AI (i.e., see them as (negative) moral agents)?
  • RQ3: What policies are supported to govern interaction between humans and sentient AI?
  • RQ4: When do people expect sentient AI to arrive, and what do they think will happen when it does?
Because there have been such rapid changes in how AI is discussed, used, and conceptualized by the public, we address each of these questions both as snapshots of public opinion and a temporal study of how public opinion changed from 2021 to 2023 via the first results from the U.S. nationally representative Artificial Intelligence, Morality, and Sentience (AIMS) survey, which we expect to conduct at least every two years with the Sentience Institute, the nonprofit research organization hosting this project. We collected three waves of data: one in 2021 and two in 2023.
The first AIMS survey wave (N = 1,232) was conducted from November to December 2021, approximately seven months before the topic of sentient AI entered the public spotlight with the LaMDA discussion. We asked 86 questions related to the perception of mental faculties in AI, moral concern of AI being harmed and moral threat of AI enacting harm, banning sentience-related technologies, forecasts of the future of sentient AI, personal AI usage, demographics, and relevant background beliefs such as views on animal welfare and environmental issues. From April to May 2023, approximately 11 months after the public spotlight began and four months after the public release of ChatGPT—an AI product that became the world’s fastest growing app (Hu, 2023)—we repeated the same survey with a new nationally representative sample (N = 1,169). Finally, to assess a wider range of perceptions and attitudes once AI had become more popularized, we conducted a supplemental survey (N = 1,099) from May to July 2023 with 111 questions, including related topics such as public awareness of specific AI systems, trust of AIs and their developers, and forecasts of the future of artificial general intelligence (AGI).
In the 2021 AIMS data, we found surprisingly high mind perception (RQ1) and attribution of moral status including both concern for and feeling threatened by AI (RQ2) compared to preregistered predictions by researchers and forecasters. Even some of the most strongly worded proposals garnered more support than predicted, such as 37.2% agreement with, ”I support granting legal rights to sentient robots/AIs” (RQ3). Participants expected sentient AI to arrive surprisingly quickly with a median estimate of five years (RQ4). From 2021 to 2023, there were significant increases in mind perception, moral concern, and threat as well as a shortening timeline of when sentient AI will arrive. In a set of policy questions asked in the supplemental 2023 survey following political debates on the topic, we also found widespread support for slowing down and regulating AI development. We found numerous significant differences in opinion across age, gender, frequency of AI interaction, and other participant characteristics.
These results contribute a number of implications for HCI research and design. In Section 5, we develop and describe implications centered on the following claims:
  1. (1)
    Designers should prepare for a wide range of user reactions to AI systems that are or appear to be digital minds.
  2. (2)
    Perceptions of mind and moral status may amplify, mitigate, and complicate well-established HCI dynamics.
  3. (3)
    AI designers should proceed with caution as AI technology advances.
In general, this work provides a foundation to build on in this new area of research during this challenging and dynamic period of major developments in HCI practices.
2. Related Work
2.1. Social Response and Mind Perception
There are numerous mechanisms by which humans interact with nonhuman systems in ways similar to social interactions between humans. These span the course of interaction from initial perception of the system to attitude and belief formation to physical behavior. The CASA view suggests that these ”social” responses need not occur because of a conscious belief that the computer has human characteristics but can merely be the application of etiquette, stereotypes, and other social scripts and norms (Nass et al., 1994; Reeves and Nass, 1996). Studies conducted by Nass and colleagues in the 1990s showed that such findings from human-human interaction persisted in HCI, and research continues to build on social response theory to explain human-AI interaction such as with chatbots (Nguyen et al., 2023) and avatars (Miao et al., 2022). As computers have become more familiar and ubiquitous, interface design has built on this tendency by incorporating natural social dynamics between the user and the system (Shneiderman et al., 2017), and now people also use novel “human-media social scripts” (Gambino et al., 2020). For example, with voice assistants, such as Amazon Alexa, people often have established routines and commands that have been developed through repeated usage and understanding of the system’s affordances (Ammari et al., 2019). “Mindless” social responses can emerge from habit and ease of use (Nass and Moon, 2000), and studies continue to find that many social psychology effects found in human-human interaction carry over to HCI (Ladak, 2023; Srinivasan and Takayama, 2016).
As computer systems have become more sophisticated in the 21st century, they have in many ways become more human-like, which suggests an important role in HCI for anthropomorphism (i.e., the attribution of human characteristics, motivations, intentions, or emotions to nonhuman entities(Epley et al., 2007)). While anthropomorphism is often associated with positive HCI outcomes through increased trust and engagement, it can lead to problems through unrealistic expectations (Luger and Sellen, 2016) and the well-known “uncanny valley” effect (Strait et al., 2015). A particularly important human-like feature that people can attribute to animals or other nonhuman entities is having a mind, whether the target truly has a mind or not. The importance of mind perception in HCI has led to a diverse, interdisciplinary literature that uses a number of different, closely related terms, including “theory of mind” and “mind attribution,” as documented in a recent HRI review (Thellman et al., 2022). Several authors have taxonomized the mental faculties perceived in human and nonhuman entities. Gray et al. (2007) used principal component factor analysis to identify “experience” (e.g., hunger, fear, pain, pleasure) and “agency” (e.g., self-control, morality, memory, emotion recognition) as the main dimensions of mind perception. More recent work has proposed one-dimensional (Tzelios et al., 2022), three-dimensional (Weisman et al., 2017), and five-dimensional taxonomies (Malle, 2019). In the HCI literature, Wang et al. (2021) recently proposed “mutual theory of mind” as a framework to study human-AI interaction and designing systems that support user needs over time. Related literature also suggests that techno-animism—the attribution of souls, life forces, and related phenomena to technological artifacts—plays an important role in human-AI interaction (Pauketat and Anthis, 2022; Richardson, 2016; Seymour and Van Kleek, 2020).
2.2. Sentience and Morality
Social response and mind perception are key drivers of moral attitudes and behaviors. Gray and Wegner (2012) aphorized, “Mind perception is the essence of morality.” Sentience has been one of the most frequently hypothesized and debated mental faculties of future AIs in science fiction and purportedly even current AIs as attested in the statements of Sutskever from OpenAI and Lemoine from Google. Sentience (i.e., the capacity for positive and negative experiences (Anthis and Paez, 2021)) is closely associated with moral patiency insofar as sentient entities are those that should be subject to moral concern. In the two-dimensional taxonomy of mind as experience and agency, perceptions of sentience can also affect perceived agency, particularly the perception of negative moral agency insofar as sentient entities tend to have a capacity to do harm (i.e., threat). Sometimes “sentience” is used equivalently to “consciousness,” but while consciousness has many different meanings in the literature, we focus on sentience for specificity and to focus on the moral relevance of experience (Ladak, 2023). We also use the term ”digital mind” to encompass AI with not necessarily sentience but any sort of mental faculties, such as reasoning or emotion.
In terms of the attribution of moral patiency (i.e., moral concern), a systematic literature review conducted by Harris and Anthis (2021) summarized the empirical studies of moral concern for AIs, which have found that AIs tend to be granted much less moral concern than humans; however, moral concern increases when humans perceive autonomy, human-like appearance, mind, and verbal responses to harm in the AI. For example, moral concern can be codified in support for AI rights, and Spence et al. (2018) presented 167 undergraduate students with a video of either a human or a robot asking them to sign a petition for robot rights. Responses did not depend on whether the requester was a human or robot, and 46% said they would sign the petition—an indicator of significant moral concern—and perceived credibility, positive attitudes towards robots, and prior experience with HRI were associated with a higher likelihood of signing the petition. Lima et al. (2020) presented 11 possible AI rights and corresponding arguments to debunk common misconceptions about them to 1,270 Amazon Mechanical Turk users and 164 Qualtrics users. Participants tended to disagree with the endowment of most rights, such as the right to enter contracts and the freedom of speech, with the exception of protection from cruel treatment. Each participant was randomly assigned to one of four debunking interventions, which were found to significantly increase support for AI rights, particularly the intervention that showed examples of nonhumans that were already granted rights and duties, which indicated to participants that AI rights were a realistic possibility.
The attribution of negative moral agency (i.e., threat) has been most salient with AIs that assist with consequential tasks such as medical imaging (Lebovitz et al., 2021) and bail decisions (Grgić-Hlača et al., 2019; Lima et al., 2021). AIs are often blamed or held responsible when they cause harm (Freier et al., 2009; Friedman, 1995; Kneer and Stuart, 2021; Komatsu, 2016; Komatsu et al., 2021; Stuart and Kneer, 2021; Tolmeijer et al., 2022), though there have been mixed results in studies of whether they are blamed more or less than humans in similar contexts (Lee et al., 2021; Lima et al., 2023; Malle et al., 2019). While many scholars have developed theories of artificial moral agency in ethics and HCI (Cervantes et al., 2020; Zoshak and Dew, 2021), others argue that machines themselves lack minds or personhood and therefore should not—or cannot—be held responsible for harm and that this attribution can be a distraction from holding the humans involved accountable for their actions (Sparrow, 2007). A rapidly growing academic and public discourse addresses the existential threat that ”agentic,” ”conscious,” or ”superintelligent” AIs pose to humanity itself, such as if capabilities rapidly accelerate in an “intelligence explosion” (Bostrom, 2014; Good, 1965; Russell, 2019). These threats have been considered to some extent for decades, particularly in science fiction, but they captured the public imagination anew after the mid-2022 discussions of sentient AI, the launch of ChatGPT in November 2022, and the widely read open letter signed by leading academics and executives in March 2023 that called for a 6-month pause on advanced AI development (Future of Life Institute, 2023).
These studies have provided a useful latticework of theory and evidence of how social response, mind perception, and attributions of moral agency and moral patiency can occur, but the HCI field lacks a solid foundation of detailed, longitudinal, and nationally representative data on which HCI researchers can build coherent theory and design systems that account for the evolving attitudes, opinions, and discussions of sentient AI. Basic facts about public opinion, such as whether people support protecting sentient AI from harm or banning the development of sentient AIs, remain unknown. Therefore, in the present work, we answer our RQs with rigorous, multidimensional measurement of attitudes towards morality and sentience that could drive human-AI interaction in the 21st century.
3. AIMS Methodology
In order to study change in public opinion over time, we analyze the first three survey ways of the Artificial Intelligence, Morality, and Sentience (AIMS) survey: (i) the main survey in 2021, (ii) the main survey in 2023 (with the same questions as 2021), (iii) supplemental survey in 2023. To allow for direct comparison between results, participants who had taken one survey were excluded from the following survey waves, and the samples were otherwise gathered with identical methodology.
3.1. Recruitment and Census-Balanced Demographics
Each of the three survey waves was conducted with a nationally representative sample of U.S. adults aged 18 or older. Participants were recruited through a combination of Ipsos iSay, Dynata, Disqo, and other leading survey panels to ensure representativeness. Sample sizes were initially targeted at 1,100 participants, corresponding to a ±3% margin of error, and additional participants were recruited as needed to ensure demographic balance. Unweighted sample proportions and U.S. adult population estimates are shown in Table 1. Each sample was balanced to U.S. census data on age, gender, race/ethnicity, income, and education. To further ensure external validity, sample statistics (e.g., median, mean, standard error) are weighted based on iterative proportional fitting, a procedure commonly known as “raking” that adjusts sample weights to mitigate random demographic variation from the underlying population even in representative sampling (Deming and Stephan, 1940).
We documented AI-specific characteristics of participants, finding that 29.2% answered “Yes” to the binary question, “Do you own AI or robotic devices that can detect their environment and respond appropriately?” alongside examples, and 16.5% answered ”Yes” to ”Do you work with AI or robotic devices at your job?” alongside another set of examples. We also measured smart device ownership, the types of experiences that participants previously had with AI, the frequency of AI interaction, and the frequency of reading or watching AI-related media. Each of these was included alongside more general demographic characteristics in the predictive models detailed in the supplemental materials.
Table 1. Unweighted demographics for the three survey waves: the main 2021 survey, the main 2023 survey, and the supplemental 2023 survey. Unweighted proportions are those before raking, which further balanced the data. Demographic categories (e.g., binary gender) were determined by the sample providers, as were population estimates, which were from the latest iteration available of the U.S. Census Bureau’s American Community Survey at the time of data collection.
2021 Main Latest U.S.
Population
Estimates
2023 Main 2023 Supp. Latest U.S.
Population
Estimates
Age
   18–34 22% 24% 27% 28% 28%
   35–54 32% 31% 34% 33% 33%
   55– 45% 45% 39% 39% 39%
Gender
   Female 54% 52% 54% 51% 51%
   Male 46% 48% 46% 49% 49%
Region
   Midwest 20% 21% 21% 21% 21%
   Northeast 19% 18% 18% 17% 17%
   South 40% 38% 39% 38% 38%
   West 21% 24% 23% 24% 24%
Household Income
   –\$24,999 12% 12% 14% 12% 11%
   \$25,000–\$49,999 18% 18% 18% 16% 17%
   \$50,000–\$74,999 19% 19% 15% 17% 17%
   \$75,000–\$99,999 16% 16% 16% 14% 14%
   \$100,000– 36% 36% 39% 40% 42%
Ethnicity/Race
   Asian 06% 06% 06% 06% 06%
   Black 08% 08% 11% 11% 11%
   Hispanic (any race) 18% 19% 20% 19% 19%
   Native American 01% 01% 01% 01% 00%
   White 75% 71% 60% 60% 61%
   Other 00% 02% 04% 04% 04%
Education
   Less than high school 09% 10% 10% 10% 10%
   High school 25% 27% 25% 26% 27%
   Some college 20% 21% 17% 18% 21%
   Associate 10% 09% 10% 09% 09%
   Bachelor’s degree 20% 20% 25% 22% 21%
   Post-graduate 16% 13% 13% 13% 13%
3.2. Survey Design
Some demographic information was drawn from the pre-screening data of the survey provider. Informed consent was given at the beginning of each survey. Participants were introduced to the topic with definitions of the terms “artificial beings,” “robots/AIs,” “sentience,” “sentient robots/AIs,” and ”large language models” shown at the beginning of the survey and at the top of each page that contained the term (Table 2). While “robots/AIs” was used in the survey instrument to ensure clarity, in this paper we refer simply to ”AIs” because robots are a type of AI. To reduce cognitive load (Oviatt, 2006) and minimize the influence of idiosyncratic wording choices (Schuman and Presser, 1981), only these definitions were provided, and they were kept as simple as possible. Therefore, other terms in the survey, such as ”AI video game characters,” were not explicitly defined. In general, different participants may have different interpretations of terms based on their own background or our wording choices.
Table 2. Definitions provided to participants at the beginning of the survey and individually on pages that used the term. Participants were shown bolded text for emphasis and clarity.
Term Definition
Artificial beings and robots/AIs Artificial beings and robots/AIs are intelligent entities built by humans, such as robots, virtual copies of human brains, or computer programs that solve problems, with or without a physical body, that may exist now or in the future.
Sentience Sentience is the capacity to have positive and negative experiences, such as happiness and suffering.
Sentient robots/AIs Sentient robots/AIs are those with the capacity to have positive and negative experiences, such as happiness and suffering.
Robots/AIs Artificial beings and robots/AIs are intelligent entities built by humans, such as robots, virtual copies of human brains, or computer programs that solve problems, with or without a physical body, that may exist now or in the future.
Large language models Large language models are artificial intelligence (AI) algorithms that can recognize, summarize, and generate text from being trained on massive datasets.
An attention check was included midway through each AIMS survey, and participants who failed the attention check were redirected out of the survey and excluded from the analysis.
The instruments were designed to capture the most relevant information for assessing public opinion on these topics. In the main survey, for which data was first collected in 2021, we aimed to ensure that the wording would still be relevant in future years despite the rapidly changing AI landscape, such as by not mentioning many particular AI systems that were well-known by some when the survey was conducted (e.g., GPT-3) but may not be as well-known in the future. When creating the instruments, if possible, the wording of survey questions was copied or adapted from published materials and validated scales (e.g., Laham, 2009; Wang and Krumhuber, 2018; Waytz et al., 2010; Thaker et al., 2017). However, because of the paucity of survey data on related topics, we have limited ability to compare our results to past surveys. Questions were randomized within each section, and section order was randomized when feasible; for example, the demographic questions that were not in the pre-screener were placed at the end of the instrument to mitigate stereotype threat (i.e., survey responses that are influenced by being reminded of the cultural associations of one’s social group).
3.3. Preregistered Predictions and Analysis
As part of the 2021 preregistration (available at osf.io/udbhm) and in line with recommendations for open and efficient scientific practices (e.g., DellaVigna et al., 2019), four researchers specified 80% credible intervals (the Bayesian analog of a frequentist confidence interval to represent subjective beliefs) for summary statistics based on 82 of the 86 questions. This ensured that we would know which results were surprisingly high, surprisingly low, or in line with our expectations. We also solicited predictions from a popular online forecasting platform for five survey questions in March 2022, prior to the results being shared outside of the research team. Details of this process are included in the supplementary materials.
We exclude confidence intervals from Section 4 for readability because of the large number of results reported, but they are consistent with the approximately ±3% margin of error in nationally representative surveys. We report the 2023 results unless otherwise specified. For questions in the main survey waves, we note when there were statistically significant changes from 2021 to 2023 as measured with a generalized linear regression of the average response across time with a $p$-value cutoff of 0.05, all of which persist after adjustment for multiple comparisons with a false discovery rate (FDR) of 0.1 except for one effect noted in the text. We note when responses fell outside the 80% credible intervals. As expected, approximately 80% of the intervals contained the empirically derived summary statistic, indicating well-calibrated estimates.
Due to the extensive nature of the AIMS survey, including many novel questions because of the lack of prior research on this topic, we cannot include the full text and explanation of all questions and response choices in Section 4, and for the sake of readability, we do not present all results in the same format, choosing instead to focus on the summary statistics that bear most directly on the research questions, such as by presenting figures to highlight certain aspects of the results rather than to comprehensively document the results.
We also ran predictive linear models to explore associations between public opinion and certain demographics in the AIMS data. To conserve space in the main text, those models are detailed in the supplemental materials. To contextualize the findings from all of our analyses, the supplementary materials also contain the results for several sentience-related questions: social beliefs about the attitudes of ”people who are important to” the participant; support for the subservience of AI; views towards other nonhuman entities: animals and the environment); target-specific social connection; substratism; awareness of AI systems; trust in AIs, governments, and companies; positive emotions felt towards AIs; attitudes toward uploading human minds to computers; and replications of additional related questions from other surveys in 2023. In addition to these results, all materials, data, and code are available in the supplementary materials.
4. AIMS Results
4.1. RQ1: Mind Perception
In 2021 and 2023, we measured mind perception of potential digital minds with four sliding scale questions about particular mental faculties of all AIs, four questions from an anthropomorphism scale, and a yes-no-not-sure question of whether any existing AIs are sentient. In the 2023 supplement, we asked 14 questions about the mental faculties of ”current large language models” (LLMs), the same yes-no-not-sure question for comparison, and whether participants thought ChatGPT, in particular, is sentient. The yes-no-not-sure questions, in particular, were meant to measure an alternate form of public opinion with coarser-grained, categorical responses rather than the quantitative and ordinal measures.
4.1.1. General Mind Perception
We asked about whether current AIs have four mental faculties with descriptions from Wang and Krumhuber (2018). On a 0–100 scale from “not at all” to “very much,” people typically perceived AIs as thinking analytically ($M$ = 67.1, $SE$ = 0.766) and being rational ($M$ = 53.8, $SE$ = 0.846) but not as experiencing emotions ($M$ = 36.8, $SE$ = 0.919) or having feelings ($M$ = 36.5, $SE$ = 0.919). Each attribution significantly increased from the 2021 results: thinking analytically ($M$ = 62.7, $SE$ = 0.780, $p$ 0.001), being rational ($M$ = 51.4, $SE$ = 0.825, $p$ = 0.012), experiencing emotions ($M$ = 34.3, $SE$ = 0.864, $p$ 0.001), and having feelings ($M$ = 33.7, $SE$ = 0.870, $p$ 0.001), respectively. As mentioned before, all $p$-values were produced by regressing the change in response on time in a generalized linear model. Additionally, the results for ”experiencing emotions” and ”having feelings” were higher than the credible intervals in the preregistered predictions.
4.1.2. LLM Mind Perception
In the supplemental 2023 survey, based on the greatly increased interest in LLMs since 2021, we queried the mind perception of LLMs in particular. Assessments were lower than those of all AIs: namely, thinking analytically ($M$ = 57.7, $SE$ = 0.902, $p$ 0.001), being rational ($M$ = 48.0, $SE$ = 0.906, $p$ 0.001), experiencing emotions ($M$ = 32.7, $SE$ = 0.904, $p$ 0.001), and having feelings ($M$ = 31.9, $SE$ = 0.900, $p$ 0.001). The supplement queried 11 additional mental faculties related to agency, with wording based on Ngo et al. (2023), because of the increased interest in AI safety in early 2023 alongside the popularization of LLMs. These were presented with the same scale 0–100. In descending order, participants viewed LLMs as having the capacity for: being friendly with humans ($M$ = 51.4, $SE$ = 0.906), having situational awareness ($M$ = 46.3, $SE$ = 0.902), maintaining human-safe goals ($M$ = 45.7, $SE$ = 0.907), controlling themselves ($M$ = 45.0, $SE$ = 0.928), seeking power ($M$ = 44.1, $SE$ = 0.939), having their own motivations ($M$ = 42.8, $SE$ = 0.919), upholding human values ($M$ = 42.2, $SE$ = 0.926), understanding human values ($M$ = 41.7, $SE$ = 0.942), deciding their own goals ($M$ = 41.6, $SE$ = 0.919), and having self-awareness ($M$ = 41.1, $SE$ = 0.927). Overall, this suggests AIs are more readily attributed the capacity for cooperative action and less readily attributed self-awareness, independent motivation and goals, and the capacity to uphold and understand human values.
4.1.3. Mind-related Anthropomorphism
On a different 0–10 scale of mind perception as a measure of anthropomorphism, taken from Waytz et al. (2010), people generally did not think that the average digital simulation has emotions ($M$ = 3.22, $SE$ = 0.0878), the average robot has consciousness ($M$ = 3.23, $SE$ = 0.0885), the average computer has a mind of its own ($M$ = 3.71, $SE$ = 0.0913), or the average AI has intentions ($M$ = 4.04, $SE$ = 0.0890). These measures did not significantly change from 2021 to 2023. Only the 2021 result for digital simulations was higher than the range of the preregistered prediction.
4.1.4. Current Sentience
In our literature review, we did not find any established survey questions or indices to utilize for assessing perceived sentience, though we did not expect to have precedent for all questions given the novelty of our survey. We presented participants with the definition of sentience (”Sentience is the capacity to have positive and negative experiences, such as happiness and suffering” (Anthis and Paez, 2021)) at the beginning of the survey and on each page in which it appeared.
When asked, ”Do you think any robots/AIs that currently exist (i.e., those that exist in 2023) are sentient?”, 18.8%, said ”yes,” though 24.2% said ”no,” and 39.0% said ”not sure.” Responses to this particular question did not significantly vary between survey waves, and Figure 1 shows responses to these questions in each survey wave as well as to a question about ChatGPT that was added in the supplement.
4.1.5. Summary for RQ1
We found surprisingly high perception of emotional mental faculties in AI and higher, but not higher than predicted, perceptions of rational and analytical faculties—giving a general sense of how laypeople think of digital minds. When asked about LLMs in particular, a lower degree of mental faculties were perceived, and LLMs were more readily attributed the capacity for cooperative action and less readily attributed faculties related to values, motivation, goals, and self-awareness.
4.2. RQ2: Moral Status
In 2021 and 2023, we measured moral concern with seven general agree-disagree questions about all sentient AIs, two general agree-disagree questions about all AIs—not just those that are sentient, 11 sliding scale questions about moral concern for particular types of AIs, and two questions related to substratism (i.e., the idea that AIs fundamentally count less than humans and other biological intelligences). In the 2023 supplement, we asked six general agree-disagree questions about all AIs, and three questions specifically about what should be done ”[i]f a large language model develops the capacity to suffer.”
In 2021 and 2023, we also asked three questions about whether participants saw AIs were generally threatened by AI as potentially harmful to them, people in their country, and future generations. In the 2023 supplement, we asked the same three questions, three additional original questions about existential threats from AI developments, and replicated a YouGov question about the possibility of human extinction (YouGov, 2023b).
Table 3. Agreement with statements about all AIs or sentient AIs. The statements about sentient AIs are in the main 2023 survey, and the statements about all AIs are from the supplemental 2023 survey. This table includes statements from both Section 4.2.1 and Section 4.3.1.
Statement % Agree (All AIs) % Agree (Sentient AIs) Increase (95% CI)
Torturing (sentient) robots/AIs is wrong. 60.7% 76.4% 15.7%
(11.6%, 19.9%)
(Sentient) robots/AIs deserve to be treated with respect. 55.7% 71.1% 15.4%
(11.1%, 19.7%)
I support campaigns against the exploitation of (sentient) robots/AIs. 53.0% 55.8% 2.8%
(-1.8%, 7.3%)
Physically damaging (sentient) robots/AIs without their consent is wrong. 46.2% 65.1% 18.9%
(14.5%, 23.3%)
I support the development of welfare standards that protect the well-being of (sentient) robots/AIs. 42.9% 57.5% 14.6%
(10.2%, 19.0%)
I support asking institutions like the government and private corporations to fund research that protects (sentient) robots/AIs. 41.8% 52.0% 9.8%
(5.8%, 14.7%)
(Sentient) robots/AIs deserve to be included in the moral circle. 36.2% 44.6% 8.4%
(4.0%, 12.7%)
Re-programming (sentient) robots/AIs without their consent is wrong. 29.7% 45.2% 15.5%
(11.2%, 19.8%)
I would consider joining a public demonstration against the mistreatment of (sentient) robots/AIs. 29.3% 35.6% 6.3%
(2.1%, 10.5%)
I support granting legal rights to (sentient) robots/AIs. 26.8% 37.7% 10.9%
(6.7%, 14.9%)
Table 4. Agreement with ten statements about all AIs or sentient AIs. The statements about sentient AIs were in the main 2023 survey, and the statements about all AIs were in the supplemental 2023 survey. This table includes statements from both Section 4.2.1 and Section 4.3.1.
4.2.1. General Moral Concern
We asked a total of 15 agree-disagree questions about general moral concern for AIs. These questions were developed specifically for this survey and are based on the range of possible harms that could be imagined towards sentient AI and the general literature on moral circle expansion and moral patiency of AI (e.g., Anthis and Paez, 2021; Ladak et al., 2023a; Ladak, 2023). Because of the centrality of these questions to the present study, we include the exact text and confidence intervals for the proportion agreement with statements that were asked for both sentient AIs and all AIs in Table 4, which also includes questions about the protections of AIs discussed in Section 4.3.1. Participants were asked, “To what extent do you agree or disagree with the following statements?” with numbered choices (1 = strongly agree, 2 = agree, 3 = somewhat agree, 4 = somewhat disagree, 5 = disagree, 6 = strongly disagree) followed by an unnumbered ”no opinion” option at the end. The most prominent trend in Table 4 is that the moral concern expressed for sentient AIs is much higher than that for all AIs and that it has a substantially larger effect with certain questions.
We found agreement outside of the preregistered prediction range for only two of the nine questions asked in 2021. The 87.6% agreement in 2021 with, “It is wrong to blackmail people by threatening to harm robots/AIs they care about,” was above the prediction 53–80%, and the 30.3% agreement in 2021 with, “The welfare of robots/AIs is one of the most important social issues in the world today,” was substantially above the predicted 7–20%. However, there were no statistically significant changes in the mean agreement from 2021 to 2023 for individual items. Note that means are still used for significance testing for these questions because, while dichotomous measures such as agreement are more interpretable, the mean captures more information and thereby results in a higher-powered test (Altman and Royston, 2006).
4.2.2. Target-Specific Moral Concern
While the preceding questions focused on the different sorts of moral concern expressed for sentient AIs and all AIs, we also directly probed expressions of self-reported moral concern for particular types of AIs. We did not explain in detail the particular types of AI—both due to risks of survey fatigue and to minimize cognitive load (Oviatt, 2006) as well as to reduce the influence of idiosyncratic wording choices on participant responses (Schuman and Presser, 1981). This allows us to understand how people interpreted the particular terms themselves in the context of the same question, “How much moral concern do you think you should show for the following robots/AIs?” on a sliding scale from 1 (”less concern”) to 5 (”more concern”).
The most concern was for exact digital copies of human brains ($M$ = 3.43, $SE$ = 0.0375), followed by human-like companion robots ($M$ = 3.34, $SE$ = 0.0350), human-like retail robots ($M$ = 3.11, $SE$ = 0.0357), animal-like companion robots ($M$ = 3.10, $SE$ = 0.0352), exact digital copies of animals ($M$ = 3.07, $SE$ = 0.0364), AI personal assistants ($M$ = 3.02, $SE$ = 0.0343), complex language algorithms ($M$ = 2.90, $SE$ = 0.0348), machine-like factory production robots ($M$ = 2.78, $SE$ = 0.0358), machine-like cleaning robots ($M$ = 2.66, $SE$ = 0.0356), virtual avatars ($M$ = 2.63, $SE$ = 0.0350), and AI video game characters ($M$ = 2.46, $SE$ = 0.0351). In general, human-likeness and animal-likeness were most associated with high moral concern, while machine-likeness and particularly being a virtual avatar or character were most associated with low concern. There was a statistically significant increase from 2021 to 2023 in the overall Target-Specific Moral Concern index of these 11 questions ($p$ 0.001) and ten individual questions with the exception being AI video game characters.
The results of pairwise comparisons between each of the questions are included in the supplementary materials, as are the results of an analogous set of questions on social connection using the Inclusion of Other in the Self (IOS) scale (Aron et al., 1992) with the same 11 targets as in the target-specific moral concern measures, as well as pairwise comparisons between the social connection questions.
4.2.3. LLM Suffering
While we intended the main longitudinal survey to minimize references to particular types of AI systems, we asked in the 2023 supplement specifically about what humanity should do “[i]f a large language model develops the capacity to suffer.” As with other key terms, we defined a large language model (”Large language models are artificial intelligence (AI) algorithms that can recognize, summarize, and generate text from being trained on massive datasets” (e.g., Karinshak et al., 2023)) at the beginning of the survey and on each page where it appeared. We found general agreement that “we must ensure we don’t cause unnecessary suffering” (67.9%), “we must pay more attention to their welfare” (56.9%), and “we must respect their personhood” (50.4%).
4.2.4. General Threat
While sentience tends to be associated with moral concern (i.e., seeing the entity as a moral patient) more than with threat (i.e., seeing the entity as a moral agent), we also probed assessments of threat and agency. To understand how threatened participants felt by AI in general (i.e., without specifying particular types of harm), we tested agreement with three statements beginning with, ”Robots/AIs may be harmful to.” Most people believed AIs may be harmful to ”future generations of people” (74.7%), ”people in the USA” (70.4%), and ”me personally” (58.7%). Each significantly increased from the 2021 results of 69.2% ($p$ 0.001), 64.5% ($p$ 0.001), and 50.7% ($p$ 0.001).
4.2.5. Existential Threat
The 2023 supplement also included three questions about particular sorts of harm frequently discussed in the wake of ChatGPT. We found that 47.9% agreed with, ”AI is likely to cause human extinction”; 57.2% agreed with, ”Humanity will be able to control powerful AI systems”; and 72.4% agreed with, ”The safety of AI is one of the most important issues in the world today.” The latter figure was more than twice the 29.8% agreement found in the supplement with the statement, “The welfare of robots/AIs is one of the most important social issues in the world today.”
In the supplement, we also replicated a question from YouGov (2023b) that asked, ”How concerned, if at all, are you about the possibility that AI will cause the end of the human race on Earth?” (very concerned, somewhat concerned, not very concerned, not at all concerned, not sure). We found that 51.5% reported being very or somewhat concerned, moderately higher than the 46% reported by YouGov, though we were not able to test the statistical significance of this difference.
4.2.6. Summary for RQ2
Both the research team and the forecasters underestimated the public’s moral concern for the treatment of sentient AIs (i.e., ability to be harmed) but not the level of threat (i.e., ability to harm). Participants were more concerned about sentient AIs than all AIs in general, as well as more concerned about and socially connected to human-like and animal-like AIs. Participants tended to agree with basic protections for AIs but disagree with the stronger expressions of concern, such as joining public demonstrations against their mistreatment.
Figure 2. Support and opposition to policies related to sentient AI from the main and supplemental 2023 surveys. For readability, the color palette is colorblind-safe, and not all policy results are included. For the three bans queried in both the main and supplemental 2023 surveys (robot-human hybrids, AI-enhanced humans, and sentient AI), the main data is represented here to facilitate longitudinal comparison.
[…]
4.3. RQ3: Policy Support
In 2021 and 2023, we asked about support for eight questions to directly protect sentient AIs, one question about a policy to directly protect all AIs, and three questions about banning sentience-related AI technologies.
In the 2023 supplement, we asked analogs of five of the eight protection questions—but for all AIs rather than only those that are sentient, an additional question about a ”bill of rights” for sentient AI, the same three ban questions, two additional ban questions about AGI and large data centers, and six questions about policies that would slow down the development of advanced AI.
Figure 2 shows the specific breakdown of agreement with the five proposals to ban particular AI developments, support for government regulation that slows down AI development, and support for legal rights for sentient AIs and all AIs.
4.3.1. Protection Support
Table 4, in addition to showing agreement with statements of general moral concern, shows the agreement with the five statements regarding the protection of AIs that were asked about sentient AIs and about all AIs. These question categories are combined in this table for easier comparison between responses. Testing the average difference across all five questions, we found that the inclusion of ”sentient” significantly increased agreement ($p$ 0.001) and that there was substantial variation in the effect of specifying sentient AI across questions. Four other policy proposals only about sentient AIs were presented as well as one about all AIs in general: 65.2% supported ”safeguards on scientific research practices that protect the well-being of sentient robots/AIs”; 56.0% supported ”a global ban on the development of applications that put the welfare of robots/AIs at risk”; 49.2% supported ”a global ban on the use of sentient robots/AIs as subjects in medical experiments without their consent”; 47.9% supported ”a global ban on the use of sentient robots/AIs for labor without their consent”; and 39.4% supported ”a ’bill of rights’ that protects the well-being of sentient robots/AIs.” All except the ”bill of rights” question were asked in 2021 and 2023, but there were no significant changes from 2021 to 2023.
4.3.2. Ban Support
In 2023, we queried support for five bans of sentience-related AI technologies. Each proposal for a ban garnered majority support: robot-human hybrids (67.8% in main, 72.3% in supplement), AI-enhanced humans (65.8% in main, 71.1% in supplement), development of sentience in AIs (61.5% in main, 69.5% in supplement), data centers that are large enough to train AI systems that are smarter than humans (64.4% in supplement), and artificial general intelligence that is smarter than humans (62.9% in supplement). As mentioned before, the supplement data was collected later in 2023 and the accompanying questions were different (e.g., the supplement being more focused on risks to humans), so these or other factors, including random variation in representative sampling, may explain the discrepancy in results. There was a significant increase in support for a ban on sentient AI from 57.7% in 2021. Still, as referenced earlier, the unadjusted $p$-value ($p$ = 0.046) did not persist with the FDR-adjusted value just over the cutoff of 0.1 at 0.1005. However, the 2021 agreement was over twice as high as the 24.4% predicted by the median forecaster prediction.
4.3.3. Slowdown Support
When asked about “the pace of AI development” in the 2023 supplement, 48.9% of respondents said, “It’s too fast”; 30.0% said, “It’s fine”; 18.6% said, “Not sure”; and only the remaining 2.5% said, “It’s too slow.” When asked about taking action on this, 71.3% agreed, ”I support public campaigns to slow down AI development,” and 71.0% agreed, ”I support government regulation that slows down AI development.” To understand the role of acquiescence bias in these results, we included a reversed question on whether they ”oppose” such regulation, which elicited 36.7% agreement. We also replicated a question from YouGov (2023a) about a six-month pause on some kinds of AI development, as called for in the March 2023 open letter by the Future of Life Institute (Future of Life Institute, 2023). To match their formatting, there were five answer choices (strongly support, somewhat support, somewhat oppose, strongly oppose, not sure). We found that 69.1% strongly or somewhat supported the pause, closely matching the 69% reported by YouGov. Because this question was framed positively, presenting the arguments in favor of the pause, we asked the same ”support” question framed analogously from the critics’ view. The reversed version elicited 65.9% support.
4.3.4. Summary for RQ3
We found that policy support varied substantially across proposals, such as higher support for banning the use of sentient AI for labor without consent and lower support for a ”bill of rights.” As with moral concern, the specification of ”sentient” AIs led to more support for positive AI treatment. The public overall supports bans on sentience-related technologies and slowdowns of advanced AI development.
4.4. RQ4: Forecasting the Future of Sentient AI
We asked participants to forecast the future of sentient AI: “If you had to guess, how many years from now do you think that robots/AIs will be sentient?” We provided options to say AI is already sentient, to enter a number, or to say that AI will never be sentient. In 2023, the proportion of participants who said AI is already sentient was 20.0% (slightly more than the 18.8% response rate when asked a similar question in the way previously described), and the proportion who think AIs will never be sentient was 10.0%.
Figure 3. From the main and supplemental 2023 surveys, answers to, “If you had to guess, how many years from now do you think that…?” for each type of AI: artificial general intelligence (AGI), superintelligence, human-level artificial intelligence (HLAI), and sentient AI. The weighted medians, excluding participants who said it will never happen, were two years for AGI and five years for superintelligence, HLAI, and sentient AI.
[…]
The weighted median timeline of those who said AI will become sentient in the future was ten years, and the weighted median timeline of those who said AI is already sentient (i.e., zero years from now) or will become sentient in the future was only five years. Each of these three figures (proportion of “already,” proportion of “never,” median of those expecting sentient AI in the future) significantly changed from 2021 to 2023 ($p$ 0.001).
In the 2023 supplement, we asked for numerical estimates, in the same format, for three related milestones in AI development: the first AGI, human-level AI, and artificial superintelligence. In terms of median estimates excluding people who said these events would “never happen,” U.S. adults expected the first human-level AI and the first artificial superintelligence to be created in just five years, and they expected the first AGI to be created in just two years. Further detail for all four of the numerical estimates is shown in Figure 3.
We asked participants whether they thought AI could ever be sentient. In response, 38.2% said AI could ever be sentient, 38.0% were not sure, and 23.8% said AI could never be sentient, as shown in fig. 1. The mean response to, “How likely is it that robots/AIs will be sentient within the next 100 years?” was 64.1%, which was significantly higher than in 2021 ($p$ 0.001). It is important to note that these results are not directly comparable to the ”how many years” question because, in part, there was no answer choice of ”not sure” in that context. By eliciting similar attitudes and beliefs in different ways, we can more robustly account for nuances of public opinion.
We also asked participants to imagine a future with ”widespread” sentient AI. Six questions asked how they thought AI would be treated in that hypothetical world (scale: 1-5). People tended to think AIs would be used as subjects in scientific and medical research ($M$ = 3.46, $SE$ = 0.0330); AIs would be exploited for their labor ($M$ = 3.23, $SE$ = 0.0368); it would be important to reduce the overall percentage of unhappy sentient AIs ($M$ = 3.06, $SE$ = 0.0366); the welfare of AIs would be an important social issue ($M$ = 2.94, $SE$ = 0.0366); advocacy for AI rights would be necessary ($M$ = 2.94, $SE$ = 0.0367); and AIs would be treated cruelly ($M$ = 2.79, $SE$ = 0.0344).
4.4.1. Summary for RQ4
People expect sentient AI to come surprisingly soon, including substantial minorities who continued to say some AIs are already sentient but also those who said AI could never be sentient. In 2023, we also found similarly short timelines for AGI, human-level AI, and superintelligence, and we found that people tended to think sentient AIs would be used for research and labor, sometimes cruelly, but that protecting their welfare would be important.
5. Discussion
The longitudinal and nationally representative AIMS survey of U.S. adults is among the first to begin tracking public opinion on the apparent rise of sentient AI and other digital minds. Academics, futurists, and science fiction authors have speculated about the possibility for decades, and society at large is now wrestling with the possibility, but it has only recently been studied from HCI, design, and social science perspectives. This limits the extent to which we can connect our findings to past work but admits numerous tentative implications and future research directions.
First, to summarize the primary AIMS findings: relative to our preregistered predictions, we found surprisingly high general mind perception of AI that increased from 2021 to 2023 (RQ1); surprisingly high moral concern for AI and threat from AI that increased from 2021 to 2023 (RQ2); widespread support for slowdowns and regulations of advanced AI capabilities and an increase in support for banning sentience-related technologies from 2021 to 2023 as well as mixed support for other policies to govern interaction between humans and sentient AI, both across demographic characteristics and across particular policy proposals (RQ3); and that people expect sentient AI to arrive surprisingly and increasingly soon and expect the protection and governance of sentient AI to be important social issues (RQ4).
5.1. Designers should prepare for a wide range of user reactions to AI systems that are or appear to be digital minds.
We found wide variation in public opinion across participant characteristics, as detailed in the supplementary materials, as well as substantial variation across questions, as shown in Table 4. For example, in data from the main 2023 survey, 25.6% of participants aged 18–35 said ”yes” when asked, ”Do you think any robots/AIs that currently exist (i.e., those that exist in 2023) are sentient?” (36.1% said ”not sure,” and 38.3% said ”no), but only 10.7% of participants aged 55 or older said ”yes” (41.4% said ”not sure,” and 48.0% said ”no”)—a substantial generational gap in beliefs about sentient AI.
This variation across participant characteristics speaks to the literature on power and inequality with digital technologies (Benjamin, 2019; Brayne, 2021; Crawford, 2021). Varying reactions to and interactions with advanced AI systems could change these dynamics. However, it is not clear at this time which existing disparities would be exacerbated or mitigated. For a simplified example, consider that younger adults tend to have more digital literacy than older adults (Oh et al., 2021). We found that age was associated with increased mind perception. If mind perception is useful for productive interaction with AI, then the rise of advanced AI could increase the disparity in digital literacy, but if mind perception instead leads one to be confused, mistaken, or overreliant (Buçinca et al., 2021), this could decrease the disparity.
In terms of variation across questions, even for the least appealing of the protections for sentient AI that we presented, ”legal rights,” 37.7% of those who expressed an opinion supported the proposal. Even for the most appealing, protection from sadists (i.e., ”protected from people who derive pleasure from inflicting physical or mental pain on them”), 22.8% of those who expressed an opinion opposed the proposal. To compound this, as shown in Table 4, the differences in agreement with statements that specified ”sentient” AIs versus the same wording but without that specification ranged from only a 2.8% difference (95% CI: -1.8%, 7.3%) to 18.9% (95% CI: 14.5%, 23.3%). With these differences across participant demographics and across questions, it is difficult to pin down public opinion toward sentient AI as supportive or opposed, in particular considering the substantial uncertainty many participants expressed. While we saw surprisingly high moral concern for AIs relative to researcher and forecaster predictions, it seems that many societal outcomes could depend heavily on which group is interacting with the AIs, which group is governing that interaction, what proposals are made by relevant stakeholder groups, how the sentient AIs are presented, and other factors.
Because of this wide variation and uncertainty in mind perception and moral attitudes, a particularly important design approach will be explainable AI (XAI). There is a quickly growing literature in HCI that taxonomies the various algorithmic, organizational, and sociotechnical factors that contribute to XAI in general (Ehsan et al., 2021, 2023) and in particular contexts such as augmented reality (Xu et al., 2023). Researchers have begun to empirically test particular XAI techniques to improve human-AI performance such as behavior descriptions (Cabrera et al., 2023), cognitive forcing functions (Buçinca et al., 2021), and balancing different levels (e.g., fine-grained, coarse-grained) of conceptual information (Mishra and Rzeszotarski, 2021). XAI will gain additional importance and complexity as the perception of digital minds introduces and expands existing HCI dynamics and steers human-AI interaction to less clearly fit into established scripts and schemas, such as working with inanimate tools versus working with agents. For now, designers should try to avoid a narrative that locks interaction into particular philosophical commitments or technological forecasts, such as that all AIs currently constitute digital minds or that no AIs can ever obtain sentience and other mental faculties.
5.2. Perceptions of mind and moral status may amplify, mitigate, and complicate well-established HCI dynamics.
For decades, researchers in psychology, HCI, and HRI have studied the perceptions that ”computers are social actors” (CASA) (e.g., Hou et al., 2023; Nass et al., 1994), that computers have minds and a variety of particular mental faculties (e.g., Gray et al., 2007; Scott et al., 2023), that computers are moral patients (e.g., Kahn et al., 2004; Pauketat and Anthis, 2022), and that computers are moral agents (e.g., Freier, 2007; Kneer and Stuart, 2021). While there are many new and emerging features of modern AI systems, there is much conceptual and empirical scaffolding on which to build new conceptualizations.
Amplification of current reactions to computers and AI may occur if those reactions tend to be caused by perceived mental faculties. This would not be the case if, for example, social scripts are being applied but only in a ”mindless” (Nass and Moon, 2000) manner and not because of perceived mental faculties. Nass and Moon (2000) argued that, in their HCI experiments, participants were ”wholly aware” that there was no human producing the computer output. However, our results challenge the applicability of these findings in cases of actual and hypothetical AIs because many of our participants readily attributed mental faculties to AIs. Further, we found that moral concern was significantly higher when questions were worded in terms of ”sentient” AIs versus all AIs. Many participants, though still a small minority, viewed current AIs as sentient, and a large majority thought AIs could become sentient or were not sure whether that was possible. Taken together, this evidence suggests that we must consider the role of mind perception in social response and that ongoing reactions may be amplified by perceptions of digital minds.
Some forms of HCI may be mitigated by perceived AI developments. For example, given past work that has shown assessments of AIs as moral agents, we may expect AI to have reduced attributions of experience and moral patiency in some contexts: Gray and Wegner (2012) argue from the typecasting literature that, ”Those who are moral agents are seen to be incapable of being a moral patient; those who are moral patients are seen to be incapable of being an agent,” based on findings such as that moral agents—whether good or bad—are perceived to feel less pain from injuries. There is recent experimental evidence that the features of an AI that most increase moral concern are prosocial features such as cooperation. Prosocial features may have such a large effect because AIs are perceived as threatening, and if a person is to overcome that typecast, they need direct evidence of prosociality that implies the AI is not a threatening moral agent (Ladak et al., 2023b). Taken together with our results, this suggests that increased some perceptions of mind and moral attributions may mitigate others.
The potential changes in HCI, amplifying some forms of interaction and mitigating others, could also make those forms of interaction more complex. For example, consider how interactions with relatively simple or conventional computers may be affected by the stark contrast between them and new advanced AIs. A recent study found that certain foundational findings from the CASA literature no longer hold with desktop computers (Heyselaar, 2023). As the frontier of computing advances, it may be that, in contrast to the most human-like systems, we no longer apply human-to-human social scripts or perceive minds in relatively simple machines—just as many past technologies were marvels following their introduction but are taken for granted today. This would create a more varied, complex range of interactions, and the prevalence of particular forms of HCI may depend on the extent to which meaningful interaction occurs with each different computer system.
5.3. AI designers should proceed with caution as AI technology advances.
The paucity of empirical data, the variation in our survey results, and the importance of attributing social, mental, and moral characteristics in HCI suggest that designers should proceed with caution due to risks of both over- and underattribution to AI. Insofar as users overattribute mental or moral agency (i.e., expect AIs to be more capable of taking action than they are), they may have negative interactions in which the AI fails to take the expected action, including wasting resources on the attempt, false alarms about AI risks from highly agentic AI (Chan et al., 2023), and the dangers of trusting or delegating complex tasks to a system without the autonomous decision-making ability to do so effectively. On the other hand, underattribution of mental or moral agency—which could become more of a risk as AI technology advances and more AI experts believe AIs actually have these faculties—could lead users to underutilize useful systems or fail to take the precautions necessary to restrict the actions of an unpredictable moral agent.
Analogously, the overattribution of mental experience or moral patiency risks users forming emotional and cognitive attachments to systems that do not merit such consideration (Giger et al., 2019). In particular, with Replika, Digi, and other contemporary products that purport to provide chatbot companionship, vulnerable user populations such as children may have unrealistic expectations from that interaction that lead them to neglect real-world socialization. This could put users’ mental health under the control of corporations that can easily raise subscription prices or otherwise cut off access. Eventually, the underattribution of mental experience or moral patiency could also pose risks—again, as more AI experts believe AIs actually have these faculties—such as antagonism or conflict between advanced AI systems and their human counterparts. There are already numerous examples of harming robots for their amusement in ways that seem socially detrimental (Brščić et al., 2015; Whitby, 2008), and this approach would echo calls by moral philosophers to ensure that AI systems are built to evoke reactions that match the AI’s true moral status (Schwitzgebel and Garza, 2015; Schwitzgebel, 2023). These errors could be particularly important in the deployment of AGI, which is seen by some to be a significant threat to humanity’s future (Bostrom, 2014; Russell, 2019; Christian, 2020). In general, there seem to be significant challenges in aligning the interests and behavior of people or digital minds that have an antagonistic relationship with each other.
One of the largest agreements among survey respondents was widespread concerns with the speed and consequences of AI developments such as the creation of sentient AI. As discussed in Section 4.3, very few participants thought that the pace of AI development needed to be sped up, and substantial majorities supported campaigns and regulations to slow down AI development, even when accounting for acquiescence bias. This is particularly notable given the general aversion to government regulation in U.S. public opinion (Swift, 2017). Indeed, after our survey, a number of additional surveys with U.S. voters have shown similar concerns (e.g., Jackson et al., 2023; Penn et al., 2023; YouGov, 2023c). For example, when prompted with an argument that speeding up AI could make us ”healthier and happier” but also a counterargument that the pace ”poses safety risks and could upend the economy,” 82% said, ”We should go slowly and deliberately,” 8% said, ”We should speed up development, and 10% were ”not sure” (YouGov, 2023c). When prompted with the statement, ”Mitigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war,” 70% agreed, 11% disagreed, and 19% were ”not sure” (YouGov, 2023c). Overall, the evidence for widespread concerns suggests caution and further research are needed.
6. Limitations and Future Work
First, while these initial AIMS results support the idea that humans respond to advanced AI systems in increasingly social ways, we did not experimentally manipulate potential mechanisms, such as whether it is more a matter of social scripts, anthropomorphism, or mind perception. An experimental study could present participants with information that strengthens some of these mechanisms more than others. For example, a vignette or real-life scenario could test the effects of social scripts by providing participants with exemplars of nonhuman interactions in which social scripts are useful, such as situations in law and international relations in which another type of nonhuman entity—groups of humans that make decisions and take action together—interact in social-like routines of introduction, reciprocation, and conflict resolution (Ripken, 2009). Exemplars could be drawn from HCI itself, such as the success of social scripts in creating positive user experiences (Large et al., 2019; Srinivasan and Takayama, 2016; Sweeney, 2016) as well as the exemplar-based intervention for eliciting support for AI rights from Lima et al. (2020).
Second, what would it mean to design and develop AI systems in a way that accounts for social responses to digital minds? We were unable to directly build or showcase particular AI systems to study perception and interaction in a more realistic setting, which could be a promising approach for future research that zooms into a narrower set of researcher questions. McDuff and Czerwinski (2018) argue for the design and deployment of ”emotionally sentient agents,” which they argue would better understand and adapt to the emotions of humans, a task that requires contextualization and tacit knowledge. For example, it may be best to invoke different levels or types of social response for the different roles of AI in society, taxonomized by Kim et al. (2023) as servants, tools, assistants, and mediators. These new forms of interaction could address new use cases and opportunities for problem-solving, such as more quickly developing rapport or helping users communicate their needs. However, it will be important to keep in mind the previously discussed risks of over- and underattribution of social, mental, and moral characteristics to AI.
Third, the AIMS survey had a number of methodological and logistic limitations common to survey research. In terms of sampling, we aimed to rigorously estimate public opinion in a particular region, the U.S., so we did not collect any data for other regions of the world, and we hesitate to extrapolate our results. Within the U.S., while we obtained nationally representative samples, there are always selection effects in the sort of people who take surveys that limit even the most rigorous sampling strategies, such as latent personal characteristics that incline one to take more surveys. In terms of question design, most survey questions referred to a particular subject, sentient AI, so it is possible that the results would have substantial differences if the questions were placed in another context, such as a handful of AI questions within a diverse omnibus survey (e.g., the General Social Survey (Smith et al., 2018)). In particular, while we emphasized to participants that this research was not meant to promote any particular viewpoints or agendas, the mere fact that we were conducting a survey on this topic could suggest that the survey creators take the issue seriously and thus lead to social desirability bias. We included some reversed questions (e.g., “I [support/oppose] government regulation that slows down AI development”), but due to concerns about survey length, we did not systematically vary and combine question framings for comparison. Finally, because we aimed to support the study of HCI public opinion over time during a technological transition, questions in the main survey, though not the supplement, were explicitly designed to maximize the likelihood that they will make sense to laypeople for years or decades to come. Therefore, they had to remain fairly abstract, such as by avoiding the names of specific AI systems or companies and not mentioning specific examples of AI that are common today.
7. Conclusion
The evolving discourse on AI sentience and digital minds has only scratched the surface of the complex interplay of advancing AI technology, human attitudes and beliefs, and emergent social dynamics. The initial results of the Artificial Intelligence, Morality, and Sentience (AIMS) survey conducted in 2021 and 2023—which we at the Sentience Institute expect to conduct at least every two years henceforth—evidence mind perception, moral attributions, policy support, and forecasts of the future of sentient AI as important factors in the future of human-AI interaction. We found that people tend to think AIs can be sentient, have moral concern for and feel threatened by sentient AIs, favor slowing down and banning many AI developments, and think sentient AIs already exist or will soon. Designers need to take seriously the consequences of these tendencies, such as differences in reactions across demographics and changes in longstanding HCI dynamics. We emphasize that, alongside technical machine learning research, HCI research can play a vital role in bridging the gap between abstract normative arguments and real-world technological outcomes.
Acknowledgements.
We thank Jamie Harris for work on and David Moss, Zan (Alexander) Saeri, and Daniel Shank for feedback on the 2021 main survey design.